If 500g of marble are decomposed completely and knowing that it contains 90% calcium carbonate, calculate the mass of calcium oxide and the volume of carbon dioxide ( at S.T.P.) produced.
1
Expert's answer
2020-05-19T08:59:14-0400

The decomposition of calcium carbonate consists in a following reaction:

CaCO3 "\rightarrow" CaO + CO2"\uparrow" .

According to this reaction, the number of the moles of calcium oxide, carbon oxide and calcium carbonate relate as:

"n(CaCO_3)=n(CO_2)=n(CaO)" .

The number of the moles of calcium oxide is its mass (500 g · 0.9 = 450 g) divided by its molar mass (100.09 g/mol):

"n(CaCO_3) = \frac{450 \text{ g}}{100.09 \text{ g/mol}} = 4.50" mol.

Therefore, the number of the moles of CO2 and of CaO are:

"n(CaO)=n(CO_2) = 4.50" mol.

The mass of calcium oxide produced is a product of its number of the moles and its molar mass (56.08 g/mol):

"n(CaO)=4.50\u00b756.08 = 252" g.

The volume of carbon dioxide produced can be calculated from the ideal gas law:

"V(CO_2) = \frac{nRT}{p} = \frac{4.50\text{ mol}·8.314 \text{ m}^3\text{Pa K}^{-1} \text{mol}^{-1}·273.15\text{ K}}{10^5\text{ Pa}}"

"V(CO_2) = 0.102" m3.

Answer: If 500g of marble are decomposed completely and knowing that it contains 90% calcium carbonate, 252 g of calcium oxide and 0.102 m3 of carbon dioxide ( at S.T.P.) produced.